The length of a rectangle is 8 inches greater than its width. Find the width of the rectangle if its area is 105 square inches.
Thank you

Answer by fractalier(6550)

Let the width be w. The length is then w+8. We multiply them to get area.
A = LW
A = (w+8)w
105 = w(w+8)
w^2 + 8w - 105 = 0
(w + 15)(w - 7) = 0
w = -15 (impossible) or
w = 7 inches